OpenStudy (anonymous):

Given two circles whose radii are 35" and 28", the ratio of their areas is 10 to 16 5 to 4 25 to 16 25 to 8 Thanks for the help, will fan & medal!

OpenStudy (phi):

You could do it the "hard way" Area of a circle is \[ A = \pi r^2 \] so the ratio of areas is \[ \frac{\pi \cdot 35\cdot 35}{\pi \cdot 28\cdot 28} \] which you can simplify but in general, similar shapes with *sides* in the ratio of " a : b" will have their areas in the ratio of a^2 : b^2

OpenStudy (phi):

so simplify I would write 35 as 5*7 and 28 as 4*7 , so the ratio of areas is \[ \frac{ \pi \cdot 5 \cdot 7 \cdot 5 \cdot 7 }{\pi \cdot 4 \cdot 7 \cdot 4 \cdot 7 }\] now cancel things that are in both the top and bottom \[ \frac{ \cancel{\pi} \cdot 5 \cdot \cancel{7} \cdot 5 \cdot \cancel{7} }{\cancel{\pi} \cdot 4 \cdot \cancel{7} \cdot 4 \cdot \cancel{7} }\]